OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form

OpenStudy (ash2326):

Standard for of a equation \[y=mx+c\] m=slope c=intercept on y-axis here slope is given as 2, so m=2 \[y=2x+c\] The line passes through (-1,5 ) so y=5 and x=-1 \[5=2(-1)+c\] we get \[5=-2+c\] add 2 to both sides, we get \[c=7\] so \[y=2x+7\] in standard form \[2x-y=-7\]

hero (hero):

I know what standard form is directrix (rolls eyes). I wasn't done posting, but okay.